OpenStudy (ujjwal):

\[\lim_{n\to\infty }\sqrt[n]{1^2+2^2+...+n^2}\]

OpenStudy (anonymous):

i have a guess maybe write the closed form of \[\sum_{k=1}^n k^2\]

OpenStudy (anonymous):

\[\lim_{n\to \infty}(\frac{n(n+1)(2n+1)}{6})^{\frac{1}{n}}\]might be a good start

OpenStudy (anonymous):

then perhaps l'hopital

OpenStudy (anonymous):

just a guess mind you, but it seem possible now take the log etc

OpenStudy (ujjwal):

I wrote it that way. what i am confused with is that do i need to show individually that n1/n, \((n+1)^{1/n}\) and other terms converge to 1? Basically here all terms will converge to 1. And additionally, i haven't learnt l'hopital so far. So, i am not supposed to use that. I guess. I was wondering if there is a better or a shortcut way.

OpenStudy (anonymous):

\[\begin{align*}\lim_{n\to \infty}\left(\frac{n(n+1)(2n+1)}{6}\right)^{1/n} &=\left(\lim_{n\to\infty}\frac{1}{6^{1/n}}\right)\left(\lim_{n\to \infty}n^{1/n}\right)\\&\quad\times\left(\lim_{n\to\infty}\left(n+1\right)^{1/n}\right)\left(\lim_{n\to\infty}(2n+1)^{1/n}\right) \end{align*}\] The first limit clearly converges to 1. The second limit can be evaluated as follows: \[\begin{align*} \lim_{n\to\infty}n^{1/n}&=\lim_{n\to\infty}\exp\left(\ln n^{1/n}\right)\\\\ &=\lim_{n\to\infty}\exp\left(\frac{1}{n}\ln n\right)\\\\ &=\exp\left(\lim_{n\to\infty}\frac{\ln n}{n}\right)\end{align*}\] which can be evaluated using the squeeze theorem (since you're l'Hopital-less).

OpenStudy (anonymous):

Try using the same reasoning for the remaining limits.